OpenStudy (anonymous):

a quality y is initially -68 and decreases at a rate of 18 per day. find an equation y=mx+b that models the quality y after x units of time. what is the equation that models the given situation?

OpenStudy (goformit100):

What's your opinion for this question ?

OpenStudy (anonymous):

its a equation

OpenStudy (goformit100):

Show your working

OpenStudy (anonymous):

-68x18

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

so we know that y=mx+b so think about how your start point is -68 that tells you that your initial y is -68 so which part of y=mx+b tell you your initial y value

OpenStudy (anonymous):

18

OpenStudy (anonymous):

which part of y=mx+b changes the y value is it the m or the b since you know the arent function is y=x

OpenStudy (anonymous):

parent*

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ok so since b changes the y value and the problem told you your initial y is -68 where should -68 go in y=mx+b just retype y=mx+b with where you think -68 should go

OpenStudy (anonymous):

y=mx-68

OpenStudy (anonymous):

ok good. Now we have to find m which we say tells us the slope of the graph. What does slope mean to you

OpenStudy (anonymous):

y=18x-68

OpenStudy (anonymous):

m=18, b=-68 where m is the slopeGlossary slope The steepness of a line expressed as a ratio, using any two points on the line. and b is the y-interceptGlossary y-intercept The value of y at the point where a curve crosses the y-axis.

OpenStudy (anonymous):

yes that is correct but the idea is that the problem gives you a number and tells you it changes at a constant rate. Since it changes at a constant rate it means that it is a linear graph that is how you identify the graph type. Then you take the initial amount given and replace the b since we know linear equations are y=mx+b. After that you know your slope is that change so you take that constant rate and get m so you complete y=mx+b

OpenStudy (anonymous):

yeah you got it that's good